OpenStudy (anonymous):

If a dictionary is sold at its marked price, then the percentage profit is 30%. If the dictionary is sold at a 20% discount on its marked price, then the profit is $5. Find the cost of the dictionary.

OpenStudy (anonymous):

what was the ans to prev q you posted?

OpenStudy (anonymous):

$120

OpenStudy (anonymous):

so i was right..yay :)

OpenStudy (anonymous):

this one is more tricky though

OpenStudy (anonymous):

@dlearner and yet i cannot get the answer

OpenStudy (anonymous):

the previous question

OpenStudy (anonymous):

did you see my post?

OpenStudy (anonymous):

yes, but i still don't understand

OpenStudy (anonymous):

wait a min

OpenStudy (anonymous):

i'll post in that question

OpenStudy (anonymous):

ok thankd

OpenStudy (anonymous):

s*

OpenStudy (anonymous):

let discuss this question

OpenStudy (anonymous):

Three variables \(m, c, d\) for marked price, cost, and discount price respectively. If a dictionary is sold at its marked price, then the percentage profit is 30%. >Profit is 30% of marked price \[ m - c = 0.30 m \]

OpenStudy (anonymous):

If the dictionary is sold at a 20% discount on its marked price \[ d = 0.20 m \] then the profit is $5. \[ d - c = 5 \]

OpenStudy (anonymous):

3 equations, three variables... shouldn't be too hard.

OpenStudy (anonymous):

Whoops, one of my equations is wrong

OpenStudy (anonymous):

If the dictionary is sold at a 20% discount on its marked price \[ d = m-0.20m \]

OpenStudy (anonymous):

Summary: \[ m-c = 0.3m \\ d-c = 5 \\ d = m - 0.20m = 0.80m \]

OpenStudy (anonymous):

i'm getting $50 for the thing

OpenStudy (anonymous):

Let \(\bf x\) denote the price of the dictionary. We know that the when the dictionary is sold at \(\bf x\) dollars, the seller gets 30% profit, i.e from the money he gets for the dictionary, 30% of that money is profit hence the profit is \(\bf 0.3x\). Now we also know that the profit when the price is decreased by 20% is $5, i.e. when the price is \(\bf 0.8x\), the profit is \(\bf $5\). But we know that from a price \(\bf x\), he makes \(\bf 0.3x\) which means that the book cost the seller about \(\bf 0.7x\). But when it's sold at 20% off, the book is getting the seller about \(\bf 0.8x\) dollars which is \(\bf 0.1x \) more than the price he bought the book at and this 0.1x profit is $5. Hence:\[\bf 0.1x=5 \implies x = $50\] @kryton1212

OpenStudy (anonymous):

how about this: see that the markoff comes directly off the profit. 30% of the initial price is profit. Taking 20% off as a discount leaves us with 70% value and 10% profit. We know this 10% profit is $5. From there we can multiply to find the answer.

OpenStudy (anonymous):

but there're only four choices which are : A. $104 B. $105 C. $125 D. $270

OpenStudy (anonymous):

wat lol then i misinterpreted the question?

OpenStudy (anonymous):

@Peter14 That's what I did too but apparently it's not right 0,o

OpenStudy (anonymous):

nah, assume the question is wrong

OpenStudy (anonymous):

nani

OpenStudy (anonymous):

what was that?

OpenStudy (anonymous):

nani = what?! "nani" is Japanese...

OpenStudy (anonymous):

it means "what?!" that's not the point...

OpenStudy (anonymous):

so now the question is, how are you going to communicate to the teacher that the actual answer is what you have and not one of the choices...

OpenStudy (anonymous):

how about if the profit is $50 instead of $5 ?

OpenStudy (anonymous):

lol?

OpenStudy (anonymous):

Then you get $500..

OpenStudy (anonymous):

i mean if the question is being like this: .......If the dictionary is sold at a 20% discount on its marked price, then the profit is $50.

OpenStudy (anonymous):

what?! $500 @@...

OpenStudy (anonymous):

yes, he's right

OpenStudy (anonymous):

ohh, but there're only that four answers for us to choose............ the question is wrong ?

OpenStudy (anonymous):

yes, as I said earlier

OpenStudy (anonymous):

oh my gosh

OpenStudy (anonymous):

this question is public exam question..

OpenStudy (anonymous):

there might be something we're overlooking...

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

$5.....

OpenStudy (anonymous):

lol @Peter14

OpenStudy (anonymous):

the answer of this question is $125

OpenStudy (anonymous):

that makes no sense at all.

OpenStudy (anonymous):

ummmmmmmmmmmmmmmmmmmm

OpenStudy (anonymous):

@kryton1212 u sure the question here is the way it is on the paper?

OpenStudy (anonymous):

yes, i am sure

OpenStudy (anonymous):

can we discuss this question after you answer my next question?

OpenStudy (anonymous):

@genius12

hero (hero):

s = selling price c = cost p = profit p = s - c \[\frac{s - c}{c} = .30\] \[.8s - c = 5\] s - c = .30c .8s - c = 5 s = 1.30c c = .8s - 5 c = .8(1.30c) - 5 c = 1.04c - 5 5 = 1.04c - c 5 = .04c 5/.04 = c 125 = c The original cost of dictionary was $125

hero (hero):

You're welcome